I know the answer to this problem is 28 however I am having trouble writing an equation for
a day care received a new sandbox.the perimeter is 84 ft and the lenght is twice as long as the width.what is the lenght of the sandbox?

Found 2 solutions by Genius, Earlsdon:
Answer by Genius(17)

a day care received a new sandbox.the perimeter is 84 ft and the lenght is twice as long as the width.what is the lenght of the sandbox?
l=2w
l=2*14
l=28
w=84/6=14 w=14
I don't know if this is what you are looking for,but I hope it at least gives you an idea.

Answer by Earlsdon(6294)

For a rectangle, the perimeter, P = 2L+2W = 84 ft.
But, L = 2W Substitute this for L in the above equation.
84 = 2(2W) + 2W Simplify and solve for W.
84 = 6W Divide both sides by 6.
14 = W
L = 2W
L = 2(14)
L = 28
The length is 28 ft.